HUBBLE SPACE TELESCOPE DAILY REPORT #2256 PERIOD COVERED: 0000Z (UTC) 11/20/98 - 0000Z (UTC) 11/23/98 Daily Status Report as of 327/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Twenty Seven Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Five Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Nine Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.4 Completed Eleven Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. The proposal completed with no reported problems. 1.5 Completed STIS/CCD/MA1/MA2 7389 (Chromospheres and Transition Layers in Hyades and Pleiades F stars) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the F class star HD28406 which is located in the Hyades star cluster. This proposal seeks to obtain additional empirical constraints on the heating mechanisms for stellar chromospheres, transition layers and coronae. A wavelenth calibration obseration was also taken. The observations were completed as planned, and no problems were reported. 1.6 Completed WF/PC-2 6664 (Stellar Ages in M 32: Exposing the Conspiracy) The WF/PC-2 was used to make observations of the galaxy NGC-221 (a.k.a., M-32) to study the stellar ages of its stars. Accurate periods and luminosities will be obtained of its RR Lyrae variable stars. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Five Sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.8 Completed Twelve Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.9 Completed WF/PC-2 6354 (Measuring Luminosity Evolution in Z = 0.3 Field Galaxies) The WF/PC-2 was used to make snapshot observations of the galactic cluster Gal-Cluster-1 located at a distance of Z = 0.3. The targets were drawn from a much larger pool of galaxies for which the internal kinematics have been measured from an extensive ground based program. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed NIC/3 7962 (NICMOS Warm-up Focus Monitoring) The Near Infrared Camera (Cameras Three) was used to make observations of the globular cluster and nebula of NGC-3603 as part of the NICMOS warm-up focus monitoring. This proposal is used to determine and monitor the optimal focus settings for NIC/2 and NIC/3 during the dewar warm-up period at the end of 1998. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/CCD/MA1 7296 (A STIS Probe of Unique ISM Lines-of-Sight Having Measured H and He Ionization Fractions) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to probe the interstellar medium around object WD2309+105 in the far ultraviolet. A wave length calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.12 Completed Five Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. As described in HSTAR 6779, a re-acquisition for the fourth iteration of this proposal failed, affecting two observations when the take data flag remained down. Otherwise, the observations were completed as planned, and no further problems were reported. 1.13 Completed STIS/MA1 7298 (A Snapshot Survey for UV bright Quasars at redshift below 3) The Space Telescope Imaging Spectrograph (MA1) was used to make ultraviolet snapshot observations of the bright quasar HE1224-081 which has a redshift (Z) below 3.0. The goal is to find quasars bright enough in the ultraviolet for spectroscopic follow-up in the range of 900-1200 Angstroms. A wave length calibration image was also taken. The observation were completed as scheduled, and no anomalies were reported. 1.14 Completed STIS/CCD/MA2 7433 (UV Spectral Dating of Galaxies) The Space Telescope Imaging Spectrograph (MA2) was used to make ultraviolet spectrographic observations of HD59984 to calculate a highly accurate library of UV stellar spectra that can be used to estimate the age and metallicity of galaxies of intermediate age {between 1 and 8 Gyrs}. The critical step in this effort is to test the model spectra against high-quality UV spectra of real stars. A wave length calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed STIS/MA1 7471 (STIS Snapshot Search for UV-bright High Redshift Quasars) The Space Telescope Imaging Spectrograph (MA1) was used to make ultraviolet observations of the bright high redshifted quasar HE0340-261. This quasar shows no signs of strong absorption in its optical spectra, and its He-II Lambda 304 A0 line is accessible with HST. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed STIS/CCD 7352 (The Nuclear Kinematics of Kinematically Distinct Cores) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the nuclear region of the kinematically distinct core of galaxy IC1459. A wave length calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.17 Completed WF/PC-2 6347 (A Search for Jets in Planetary Nebulae) The WF/PC-2 was used to make observations of the planetary nebulae (class PNe) PK032-02D1 in order to search for the presence of possible narrow jets. The current model for PNe formation indicates that jets are an unexpected phenomena, yet there is a growing body of circumstantial evidence that suggests narrow jets are common feature in PNe. The observations were completed as scheduled, and no anomalies were noted. 1.18 Completed NIC/1/2 7864 (The Host Galaxies of Radio-Quiet High- Redshift Quasars) The Near Infrared Camera (Cameras One and Two) was used to make observations of the radio-quite, high red shifted quasar MZZ11408 and its host galaxy. The underlying causes of the rapid cosmological evolution of the quasar population are a mystery around which are woven many different strands of contemporary cosmology. As described in HSTAR 6779, a re-acquisition for this proposal failed, affecting one-and-one-half hours of this six-and-one-half hour proposal and four observations when the take data flag remained down. Then, the loss of lock described in 2.1 and HSTAR 6780 occurred, possibly affecting an additional two observations. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.19 Completed Two Sets of WF/PC-2 7387 (T-Tauri Star Snapshot Survey) The WF/PC-2 was used to make observations of the stars T-Cha and FT-Tau as part of a T-Tauri class star survey. The observations were executed as scheduled, and no anomalies were noted. 1.20 Completed STIS/CCD/MA2 7334 (The Primordial Abundance of Deuterium From a Metal-Poor Damped Lyman Alpha System) The Space Telescope Imaging Spectrograph (CCD and MA1) were used to make observations of the quasar Q2206-1956 to study the primordial abundance of deuterium from a metal-poor damped Lyman-Alpha system. It should be possible to detect D absorption in many Lyman series lines and the chance of contamination by interloper H clouds is reduced by a factor of > 150. A wave length calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.21 Completed WF/PC-2 6626 (Imaging of Z>1 - Massive Star Forming Galaxies) The WF/PC-2 was used to make observations of the massive star forming galaxy SSA22E which is known to have a redshift of Z>1.0. The central regions of the galactic strips have already been imaged with HST and the current observations are to cover the samples required to finish the set. The observations were executed as scheduled, and no anomalies were noted. 1.22 Completed NIC/1/2 7319 (Composition and Heterogeneity of Io's Volcanoes From Polarimetry of their Thermal Emission) The Near Infrared Camera (Cameras One and Two) was used to make polarimetry observations of Jupiter's moon Io in order to study the composition and heterogeneity of its thermal emissions. Measurements were taken of the linear polarization of Io's 2 Mum volcanic thermal emissions while the moon was in Jupiter's shadow in order to study the lava composition and its variation among different hot spots on Io's surface. The observations were completed as scheduled, and no anomalies were noted. 1.23 Completed WF/PC-2 6458 (Imaging of QSO Host Galaxies at Z >2) The WF/PC-2 was used to make observations of the QSO/host galaxy 0033+098. QSOs of redshift larger than 2 have been resolved with ~arcsecond resolution on 4m-class ground-based telescopes. Both line and continuum flux have been detected over a few arcseconds, but the nature of the extended emission, and its relation to the nuclear energy source is unclear. The observations were executed as scheduled, and no anomalies were noted. 1.24 Completed WF/PC-2 6839 (WF/PC-2 Imagery of the Dusty H-II Region SMC N88A) The WF/PC-2 was used to make observations of nebula NEB-012409 (a.k.a. , SMC N88A) which is a H-II region inside of the Large Magellanic Cloud. This small H-II region is unusual in the metal-poor and relatively dust-free interstellar medium (ISM) environment of the SMC in that it is heavily reddened and shows a peculiar ionization structure compared to other SMC H-II regions. The observations were completed as planned, and no anomalies were reported. 1.25 Completed WF/PC-2 7501 (Expansion Parallax Distances to Planetary Nebulae) The WF/PC-2 was used to make parallax distance measurements of the planetary nebula IC1454. The largest impediment to our understanding of planetary nebulae is the lack of reliable distances. Acquired by generalized statistical methods in all but a handful of cases, individual PNe distances are often uncertain by factors of two or more. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 19 Successful : 19 Scheduled Reacquisitions : 29 Successful : 28 As documented in HSTAR 6779, the re-acquisition at 325/140319Z failed to RGA control after four unsuccessful attempts at fine lock. The following re-acquisition at 325/153731Z was successful. Proposals described in 1.12 and 1.18 were affected. There was a loss of lock at 325/164229Z following a type T3 profile slew. HSTAR 6780 was written. The proposal described in 1.18 may have been affected. 2.2 FHST Updates: Scheduled : 42 Successful : 42 2.3 Operations Notes: Using ROP SR-2, the SSR EDAC error counter was cleared twelve times. A TTR was written to describe four occasions when there were missing minor frames and data inversions on day 324. At 324/221556Z, a NICMOS single bit error occurred. HSTAR 6778 was written. The NICMOS internal log was dumped and the ground system limit was updated at 324/2313Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH